# Automatic Intermodal Loading Unit Identification using Computer Vision: A Scoping Review

### v1

@@ Abstract
The standardisation of Intermodal Loading Units (ILUs), such as containers, semi-trailers and swap bodies, has revolutionised global trade yet their efficient and robust identification remains a critical bottleneck in high-throughput ports and terminals. This paper reviews 63 empirical studies that propose computer vision (CV) based solutions. It covers the last 35 years (1990-2025), tracing the field’s evolution from early digital image processing (DIP) and traditional machine learning (ML) to the current dominance of deep learning (DL) techniques. While CV offers cost-effective alternatives for other types of identification techniques, its development is hindered by the lack of publicly available benchmarking datasets. This results in high variance for the reported results such as end-to-end accuracy ranging from 5 % to 96 %. Beyond dataset limitations, this review highlights the emerging challenges especially introduced by the shift from character-based text recognition to scene-text spotting and the integration of mobile cameras (e.g. drones, sensor equipped ground vehicles) for dynamic terminal monitoring. To advance the field, the paper calls for standardised terminology, open-access datasets, shared source code, while outlining future research directions such as contextless text recognition optimised for ISO6346 codes.
@@
intermodal loading unit identification , container code recognition , text detection , text recognition , logistics automation
journal: Transport Reviews
@@
CCA
Connected Component Analysis
MSER
Maximally Stable Extremal Regions
@@ 1 Introduction
The evolution of cargo transportation has been significantly transformed by the concept of containerisation. Before, cargo handling was labour-intensive, neither standardised nor suitable for automatisation. Handling and storage arrangements were costly and required special attention [1, 2], resulting in terminal congestion. Containerisation has facilitated automation and improved operational efficiency in cargo handling [3], while also significantly contributed to the growth of global trade [4]. The standardisation of freight handling has not only led to the shipping container but also led to a variety of standardised types of Intermodal Loading Units.
@@
The term ILU has been introduced to describe units designed to be easily transferred between different modes of transport, such as vessels, trains, and trucks, without the need to unload and reload the contents. The loading units, such as containers, interchangeable bodies (swap bodies) and semi-trailers, fall under this term [5]. These units are typically built according to specifications defined in standards such as ISO668 [6]. These specifications ensure compatibility of ILUs with different transport systems [7] for smooth movement of goods along supply chains [8]. A unique identification code, following the ISO6346 [9] standard, is assigned to each ILU for tracking purposes. This code consists of four letters, a six-digit number and a check digit, which together indicate the owner and type of the ILU.
As global trade expands, many leading ports handle tens of millions of containers annually, with some operating at or above their nominal design capacity [10]. In this high throughput, keeping track of ILUs is important to ensure efficient management [11]. The often identically looking ILUs can exclusively be distinguished by their ID codes. Therefore, a reliable system that can rapidly recognise these codes is indispensable. Standardisation and digitalisation enabled the development of various techniques for ILU identification. Among these, Computer Vision (CV)-based identification code recognition stands out as a cost-effective alternative to RFID, which achieves only about 70 % accuracy and entails high installation and maintenance costs [12, 13]. In contrast, CV-based methods can be implemented using a range of camera setups from high-end static Optical Character Recognition (OCR) gates [14], that can identify entering and exiting ILUs [15, 16], to more affordable and flexible options such as using the existing CCTV infrastructure, cameras mounted on aerial or ground vehicles, or even mobile devices, which significantly reduce the need for specialised hardware. Furthermore, CV-based identification plays a crucial role in enhancing security by providing accurate verification for customs clearance, which can support in the correct declaration of goods [17, 18, 19].
Although CV-based methods reduce human-related errors and increase the processing speed [20, 21, 22, 23] there are several challenges: Containers often arrive in rapid succession, will be stacked for storage, and can be positioned at angles that partially occlude their ID codes. Environmental factors such as rain, fog, poor lighting, or glare from sunlight can degrade image quality, while harsh conditions, such as dirt, rust, or physical damage, further complicate automated ID code recognition [14, 24]. In addition, variations in font style, colour and placement of ID codes across different operators and regions create inconsistencies that reduce the reliability of automatic recognition systems.
@@
At transfer points of the terminal to other modes of transport, each ILU must be identified and cross-checked with its associated documentation before entry or exit. When automated systems fail to achieve reliable recognition, manual verification becomes necessary, further slowing down the process. Even a brief slowdown at a busy gate can cause truck queues stretching hundreds of meters [25], disrupting yard operations and increasing turnaround times for transport vehicles. Extended waiting times contribute to congestion which increases fuel consumption and greenhouse gas emissions from idling trucks, amplifying the environmental footprint of port activities [26]. Therefore, efficient ILU identification is essential for overcoming these operational bottlenecks and sustaining the growth of containerised transport.
@@
Despite its crucial role in improving transportation and trade security, to the best of our knowledge, there are not any literature reviews on ILU identification. To fill this gap, we prepared this literature review by following Preferred Reporting Items for Systematic Reviews and Meta-Analyses extension for Scoping Reviews (PRISMA-ScR) guidelines [27] with the intention of identifying the available publications and proposed concepts.
@@
    Catalogue all empirical studies on Intermodal Loading Unit (ILU) identification using Computer Vision (CV) methods between 1990 and 2025.
@@
    Synthesise an overview of Intermodal Loading Unit (ILU) identification methods, image acquisition settings, datasets, and evaluation metrics used in these studies.
@@ 2 Terminology
This review distinguishes the roles of Digital Image Processing (DIP), Machine Learning (ML), and Deep Learning (DL) in the context of CV-based ILU identification to ensure precise use of technical terms and prevent misinterpretation across disciplines. Therefore, we provide the definitions of the general approaches:
@@
Machine Learning (ML): While DIP methods rely on fixed, hand-crafted rules and deterministic operators (e.g. edge detectors, thresholding, morphologic filtering), ML introduces adaptability by learning from data. ML methods use features engineered via DIP (such as HoG, SIFT) as input to classifiers that can discriminate among classes (e.g. character classes, object vs. non-object, scene categories). Common models include Support Vector Machines [45, 46, 47], Random Forests [48], decision trees [49] and k-Nearest Neighbors [50]. These reduce manual rule-tuning and can generalize better under variations in viewpoint, scale, and lighting, but still depend critically on quality of the features and consistency in preprocessing [51].
Deep Learning (DL): A subfield of ML using deep neural networks, such as Convolutional Neural Networks [52] or transformers [53] that learn features and classifiers jointly and end-to-end. Although DL requires substantial training data and computational resources it automatise the feature engineering step. For similar tasks depending on similar features, deep learning methods can benefit from transfer learning [54] and fine-tuning [55].
Besides the terminology for general approaches, we would like to provide definitions for CV-specific terminology:
@@
(Scene-)Text detection: Localising regions containing text in an image (e.g., street signs,) without extracting its textual content. Outputs bounding boxes (often quadrilaterals) around words or lines. Unlike character detection, it focuses on full words or lines, which makes it more robust against distractors [62].
(Scene-)Text recognition: Extracting textual information from detected text regions by converting them into machine-readable characters or words [63, 64].
@@ 2.1 Terminology Beyond OCR Gates
Initially, CV methods were applied at terminal gates, where the task is to automatically identify loading units as they entered or exited the facility. These gates are among the most vulnerable and critical points of a terminal, making reliable identification essential. The task performed there was to capture an image of the loading unit and read its identification code through DIP [44]. In the literature, this task is often referred to as Automatic Container Code Recognition (ACCR) [66, 67, 68, 69, 70, 71, 72] or transportation unit (TU) identification [73]. The focus of ACCR is on using CV to read the ISO6346 identification codes from containers at the point of entry. These entrances remained as the critical bottlenecks in terms of operational efficiency and congestion impact, so their optimisation becomes one of the key challenges in terminal management [74].
@@
However, the term ACCR is not precise. First, the word container is too narrow, since not only containers but also semi-trailers and swap bodies carry ISO6346-compliant codes (as defined in EN 13044 [75]). Second, the word recognition does not fully capture the actual tasks involved. In computer vision, recognition usually means extracting text from a clean image, especially for the mobile cameras. The term recognition refers to extraction of text from an image in the CV field. However, this presupposes that the visual input image consists solely of the target text. This assumption is rarely valid in the real-world scenarios involving surveillance cameras or mobile cameras, where text is typically embedded within complex and cluttered scenes. In these cases, an additional stage, text detection is required to localise the text regions before recognition can be performed [76]. Over time, these methods were extended beyond the gates. With advancements in computer vision and robotics, identification is now performed at quay cranes [77], yard blocks [78], and movable objects such as reach stackers [79]. These new use cases introduce more variation and complexity than the original gate-based setting.
@@
For these reasons, we propose the term Intermodal Loading Unit (ILU) identification as an inclusive designation for the task of determining which specific ILUs with ISO 6346–compliant ID codes are visible in an image. This term includes various loading units and also for the CV terminology such as ILU detection/segmentation stage introduced in recent papers [73, 80], scene-text detection and recognition. A term that covers all these stages aligned with CV terminology is important, as it would address shared challenges across the CV applications. For example, using a common terminology can link the relevant studies to the broader problems such as small object segmentation [81] or small text detection [82] which are actively studied in general purpose CV methods. Moreover, a common terminology can help in finding the distinct aspects of ILU identification. Unlike in most common text recognition needs, where a language model based encoder (e.g. TrOCR [83]) is useful due to contextual cues present in the natural language, the contextless nature of the target text in ILU identification makes this approach ineffective. Such models tend to overfit to frequently occurring text instances in training data [84], limiting their applicability in this context. Adopting the term ILU identification can provide conceptual clarity and establish a coherent basis for researchers and practitioners.
@@ 3 Methodology
As the first review on ILU identification, we applied inclusive eligibility criteria to capture the full scope of available evidence. These criteria were specified a priori before screening began. They follow the Population–Concept–Context (PCC) logic and the methodological guidance in [27]. Our review includes articles written in English, published from 1990 to 2025, that follow a scientific structure. Both peer-reviewed articles and preprints are included as long as they report quantitative results, allowing us to catalogue the latest research findings. The final selection of articles and data charting were determined through consensus among the authors, following a review of titles, abstracts, and full texts.
From the 63 eligible articles the following data items were extracted into a spreadsheet [^1] https://cloud.uni-hamburg.de/s/FHqaGLR4Ccy8ceG: journal or conference details, contributions, methods used, method class (DIP, ML, or DL), dataset specifics (availability, number of images, diversity in lighting and weather), image collection region and camera position such as fixed (CCTV) or moving (aerial or ground vehicle), evaluation metrics, results, deployment status, suggested future directions and funding. The resulting table was then analysed and discussed by the authors.
@@ 4 Results and Discussion
Table 1: Eligible article characteristics ($n=63$). Five articles involve international collaborations, which results in the count of 69 rather than the 63, in geographic regions section.
Category Characteristic Count Percentage
Year of publication 1990–1995 2 3.17 %
1996–2000 1 1.59 %
2001–2005 5 7.94 %
2006–2010 3 4.76 %
2011–2015 10 15.87 %
2016–2020 18 28.57 %
2021–2025 24 38.10 %
Geographic region Asia 55 79.71 %
Europe 8 11.59 %
North America 3 4.35 %
Africa 2 2.90%
Oceania 1 1.45 %
Funding source Not reported 32 50.79 %
Public 24 38.10 %
Public-Industry Collab. 4 6.35 %
Industry 3 4.76 %
Rank Unranked 32 50.79 %
Q1 - Scimago 8 12.70 %
A - CORE 5 7.94 %
Q2 - Scimago 5 7.94 %
B - CORE 4 6.35 %
Q3 - Scimago 4 6.35 %
Q4 - Scimago 4 6.35 %
C - CORE 1 1.59 %
Table 1 presents the characteristics of the eligible articles. The data indicate a consistent increase in publications over time, especially from 2011 onwards. This trend reflects growing interest in automatic ILU identification systems, likely driven by advancements in CV as well as greater computational power affordable to major terminals and ports.
@@
Asia contributes 79.71 % of the publications. This region handles high volumes of maritime traffic requiring efficient logistics and ILU identification systems. Countries such as China, Taiwan, South Korea, Vietnam, and Japan play significant roles in global trade, making effective logistics and supply chain management essential. Substantial investment in research and development by Asian governments and companies further supports this trend [85, 86].
@@
Public funding is reported as the main source, supporting about two out of five of the publications, indicating strong governmental backing. This support is often motivated by the potential for technological advancements to benefit national infrastructure and economic growth [87, 88, 89]. Public funding is particularly important for the logistics sector that operates with low profit margins [90, 91]. Industry funding and public-industry collaborations constitute one ninth of the publications, suggesting that while there is some industry involvement, companies largely rely on public sector investments to drive technological development. However, the presence of pure industry funding indicates a genuine market demand for companies to invest in research to remain competitive [92, 11].
To analyse the rankings of journals and conferences we followed the rankings of Scimago [^2] https://www.scimagojr.com/ for journals and CORE [^3] https://www.core.edu.au/icore-portal for conferences. Table 1 shows that more than half of the articles on automatic ILU identification task are published in unranked outlets, which is not uncommon for applied or niche technical topics [93].
@@
Among the ranked publications, Q1-Scimago and A-CORE have the highest numbers, suggesting that the topic is relevant for high-quality research venues. However, the majority of publications are spread across lower-tier or unranked venues. This distribution implies that the field could benefit from a greater emphasis on more novel contributions, increased comparability, and reproducibility of the results. This could be achieved through publicly available benchmark datasets and repositories that showcase the source code and the model weights. By adopting such research standards, ILU identification tasks with CV-based methods could be more relevant for high-ranked publication outlets.
Overall, the characteristics of the reviewed articles highlight a growing research field characterised by substantial activity in Asia and a predominance of public funding. Expanding funding sources and promoting industry collaboration may further advance automatic ILU identification through CV technologies.

### v2

@@ Abstract
Background: The standardisation of Intermodal Loading Units (ILUs), including containers, semi-trailers, and swap bodies, has transformed global trade, yet efficient and robust identification remains an operational bottleneck in ports and terminals.
Objective: To map Computer Vision (CV) methods for ILU identification, clarify terminology, summarise the evolution of proposed approaches, and highlight research gaps, future directions and their potential effects on terminal operations.
Methods: Following PRISMA-ScR, we searched Google Scholar and dblp for English-language studies with quantitative results. After dual reviewer screening, the studies were charted across methods, datasets, and evaluation metrics.
Results: 63 empirical studies on CV-based solutions for the ILU identification task, published between 1990 and 2025 were reviewed. Methodological evolution of ILU identification solutions, datasets, evaluation of the proposed methods and future research directions are summarised. A shift from static (e.g. OCR-gates) to vehicle mounted camera setups, which enables precise monitoring is observed. The reported results for end-to-end accuracy range from 5 % to 96 %.
Conclusions: We propose standardised terminology, advocate for open-access datasets, codebases and model weights to enable fair evaluation and define future work directions. The shift from static to dynamic camera settings introduces new challenges that have transformative potential for transportation and logistics. However, the lack of public benchmark datasets, open-access code, and standardised terminology hinders the advancements in this field. As for the future work, we suggest addressing the new challenges emerged from vehicle mounted cameras, exploring synthetic data generation, refining the multi-stage methods into unified end-to-end models to reduce complexity, and focusing on contextless text recognition. Addressing these challenges can result in minimised human errors, decreased turnaround times, reduced operational costs and higher resilience against disruptions.
@@
intermodal loading unit identification , container code recognition ,
journal: Transportation Research Part: C
@@ 1 Introduction
The term Intermodal Loading Unit (ILU) has been introduced to describe units designed to be easily transferred between different modes of transport, such as vessels, trains, and trucks, without the need to unload and reload the contents. The loading units, such as containers, interchangeable bodies (swap bodies) and cranable semi-trailers, fall under this term [1]. A unique identification code, following the ISO6346 [2] standard, is assigned to each ILU for tracking purposes. This code consists of four letters, a six-digit number and a check digit, which together indicate the owner, type and serial number of ILU. Identically looking ILUs can exclusively be distinguished by their ID codes.
@@
While the standardised identification of ILUs provides a framework for tracking, the scale of modern port operations introduces challenges. As global trade expands, many leading ports handle tens of millions of containers annually, with some operating at or above their nominal design capacity [3]. In this high throughput, keeping track of ILUs is important to ensure efficient management [4]. For example, at transfer points of the terminal to other modes of transport, each ILU must be identified and cross-checked with its associated documentation before entry or exit. When automated systems fail to achieve reliable ILU identification, manual verification becomes necessary, further slowing down the process. Even a brief slowdown at a busy gate can cause truck queues stretching hundreds of metres, disrupting yard operations and increasing turnaround times for transport vehicles [5, 6]. Extended waiting times contribute to congestion which increases fuel consumption and greenhouse gas emissions from idling trucks, amplifying the environmental footprint of port activities [7]. Therefore, efficient and automatic ILU identification is essential for overcoming these bottlenecks and sustaining the growth of containerised transport.
Standardisation and digitalisation enabled the development of various techniques for ILU identification. Among these, Computer Vision (CV)-based identification code recognition stands out as a cost-effective alternative to RFID, which achieves only about 70 % accuracy and entails high installation and maintenance costs [8, 9]. In contrast, CV-based methods can be implemented using a range of camera setups as shown in Figure 1. From static Optical Character Recognition (OCR) gates [10, 11], that can identify entering and exiting ILUs [12, 13], the field is shifting towards more flexible options such as camera equipped aerial or ground vehicles, or even mobile devices, which significantly reduce the need for specialised infrastructure [14, 15, 16, 17]. Furthermore, CV-based ILU identification plays a crucial role in enhancing security by providing accurate verification for customs clearance, which can support in the correct declaration of goods [18, 19, 20].
Figure 1: Schematic overview of a terminal, highlighting the use of camera-equipped aerial and ground vehicles and OCR-gates for comprehensive port monitoring.
@@
Despite its crucial role in improving port operations by reducing human-related errors and increase the processing speed [21, 22, 23, 24], to the best of our knowledge, there are not any literature reviews on CV-based ILU identification. To fill this gap, we prepared this literature review by following Preferred Reporting Items for Systematic Reviews and Meta-Analyses extension for Scoping Reviews (PRISMA-ScR) guidelines [25] with the intention of identifying the available publications and proposed concepts.
@@
    Catalogue all empirical studies on ILU identification using CV methods between 1990 and 2025.
@@
    Synthesise an overview of ILU identification methods, image acquisition settings, datasets, and evaluation metrics used in these studies.
@@ 2 Terminology
As CV-based ILU identification is an interdisciplinary field we would like to establish a common terminology to create a common ground for researchers with different background, such as robotics, artificial intelligence, and logistics. In the related literature, we realised a lack of definitions which results in ambiguity. For example, text spotting, an integrated task encompassing both text detection and text recognition, is commonly reduced to OCR, ignoring its broader scope. This ambiguous terminology complicates result interpretation, and hindering meaningful cross-study comparisons. Therefore, we provide the precise definitions for general approaches and downstream tasks.
2.1 Approaches
@@
Machine Learning (ML): While DIP methods rely on fixed, hand-crafted rules and deterministic operators (e.g. edge detectors, thresholding, morphologic filtering), ML introduces adaptability by learning from data. ML methods use features engineered via DIP (such as HoG, SIFT) as input to classifiers that can discriminate among classes (e.g. character classes, object vs. non-object, scene categories). Common models include Support Vector Machines [43], Random Forests [44], decision trees [45], k-Nearest Neighbors [46] and Reinforcement Learning [47, 48, 49]. These reduce manual rule-tuning and can generalise better under variations in viewpoint, scale, and lighting, but still depend critically on quality of the features and consistency in preprocessing.
Deep Learning (DL): A subfield of ML using deep neural networks, such as Convolutional Neural Networks [50, 51] or Transformers [52] that learn features and classifiers jointly and end-to-end. Although DL requires substantial training data and computational resources, it automates the feature engineering process. However, a major drawback of deep learning models is their black-box nature, which makes them difficult to interpret and explain [53]. For similar tasks depending on similar features, deep learning methods can benefit from transfer learning [54] and fine-tuning [55].
2.2 ILU Identification Related Computer Vision Tasks
Besides the terminology for general approaches, we would like to provide definitions for downstream tasks:
@@ 2.2 ILU Identification Related Computer Vision Tasks
Optical Character Recognition (OCR): Localising and identifying individual characters in an image. OCR is predominantly used for document text recognition and perform well under standard conditions, such as uniform backgrounds and consistent lighting (e.g., scanned documents). However, OCR techniques may fall short when dealing with text-in-wild scenarios [62], where background and lighting are highly diverse and text orientation is arbitrary. Relying on character detection in diverse scenes can increase false positives, as many objects share similar patterns with individual letters [63, 64]. Another disadvantage of character-level recognition, especially with DL methods, is that character-level recognition needs more expensive annotated data than word-level or line-level recognition due to its granularity and required labour for annotating individual characters.
For tasks involving the diverse scenes, such as identifying signs in urban areas [65], two primary subtasks are defined:
(Scene-)Text detection: Localising regions containing text in an image without extracting its textual content. Outputs quadrilaterals around words or lines. Unlike character detection, it focuses on full words or lines, which makes it more robust against distractors [66, 67, 68, 69].
(Scene-)Text recognition: Extracting textual information from detected text regions by converting them into machine-readable characters or words [70, 71, 67, 68, 69, 69].
@@ 2.3 Terminology Beyond OCR Gates
Terminal entrances remained as the critical bottlenecks in terms of operational efficiency and congestion impact, so their optimisation becomes one of the key challenges in terminal management. The earlier studies on this, applied CV methods only at terminal gates, where the task is to automatically identify loading units as they entered or exited the facility. In the literature, this task is referred to as Automatic Container Code Recognition (ACCR) [73, 74, 75, 76, 77, 78, 79] and rarely as loading unit identification or transportation unit (TU) identification [17]. The focus of ACCR is on containers at terminal gates. However, the word container does not cover all types of loading units such as semi-trailers and swap bodies that has ISO6346-compliant codes (as defined in EN 13044 [80]). Second, the word recognition does not fully capture the actual tasks involved.
@@
The term recognition refers to extraction of text from an image in the CV field. However, this presupposes that the visual input image consists solely of the target text. This assumption is rarely valid in the real-world scenarios involving surveillance cameras or mobile cameras, where text is typically embedded within complex and cluttered scenes. In these cases, an additional stage, text detection is required to localise the text regions before recognition can be performed [81]. Over time, these methods were extended beyond the gates. With advancements in CV and robotics, identification is now performed at quay cranes [82], yard blocks, and terminal vehicles such as reach stackers [83]. These new use cases introduce more variation and complexity than the original gate-based setting and therefore, text recognition is not sufficient by itself.
@@
For these reasons, we propose the term Intermodal Loading Unit Identification as an inclusive designation for the task of determining ID codes of ILUs. This term covers various loading units and also ILU detection/segmentation stage prior to text detection and recognition steps, introduced in some studies [17, 84]. A term that covers all these stages aligned with CV terminology is important, as it would address shared challenges across the CV applications. Using a common terminology can link the relevant studies to the broader problems. For example, different scales of objects and text is a problem that was introduced with the mobile cameras for ILU identification task. However, small object segmentation [85] or small text detection [86] havebeen actively studied in general purpose CV methods. Moreover, a common terminology can help in finding the distinct aspects of ILU identification from the broader tasks. Unlike in common text recognition, where a language model based encoder (e.g. TrOCR [87]) is useful due to contextual cues present in the natural language, the contextless nature of ISO6346-compliant ID codes makes this approach ineffective. Such models tend to overfit to frequently occurring text instances in training data [15], limiting their applicability in this context. Adopting the term ILU identification can provide conceptual clarity and establish a coherent basis for researchers and practitioners.
Having established the terminology and challenges, we now turn to the methodology used for systematically review the literature.
@@ 3 Methodology
This literature review was conducted in accordance with the PRISMA-ScR guidelines [25] and eligibility criteria were specified a priori and piloted before the screening began. These criteria follow the Population–Concept–Context (PCC) logic and Table 1 details the inclusion and exclusion rules together with their rationales.
Table 1: Inclusion and exclusion criteria for systematic review of ILU code recognition.
Domain Inclusion rule Exclusion rule
Date range 1 Jan 1990–31 August 2025 << 1990 or after cut-off
Language English full text Non-English; abstract-only
Publication type Peer-reviewed journal / conference paper; pre-print (quantitative) Patents, standard, theses, news or editorials
Population (P) ILUs (containers, swap-bodies, semi-trailers) Other objects unless ILUs also analysed
Concept (C) CV localisation and recognition of alphanumeric ILU code RFID/barcode solutions
Context (C) Fixed or mobile cameras in ports, terminals, rail yards Depots
Reporting completeness Method detail plus numeric outcomes (precision, recall, F1, mAP, read-rate) No quantitative results
Duplicates / accessibility First complete, accessible full text Duplicates; inaccessible full text
As the first review on ILU identification, we applied inclusive eligibility criteria to capture the full scope of available evidence. Our review includes articles written in English, published from 1990 to 2025, that follow a scientific structure. Both peer-reviewed articles and preprints are included as long as they report quantitative results, allowing us to catalogue the latest research findings. The final selection of articles and data charting were determined through consensus among the authors, following a review of titles, abstracts, and full texts.
@@
From the 63 eligible articles the following data items were extracted into a spreadsheet [^1] Available in supplementary materials.: journal or conference details, contributions, methods used, method class (DIP, ML, or DL), dataset specifics (availability, number of images, diversity in lighting and weather), image collection region and camera position such as fixed (CCTV) or mobile (aerial or ground vehicle), evaluation metrics, results, deployment status, suggested future directions and funding. The resulting table was then analysed and discussed by the authors.
@@ 4 Results and Discussion
Figure 2: Geographic distribution of articles by country (n=63). Five articles involve international collaborations, which results in 69 for country mentions rather than 63.
@@
Most of the published articles were written by the authors with Asian affiliations (79,71 %) as shown in Figure 2. This is related with the high volumes of ILU handling in Asian ports and terminals, which creates a need for efficient ILU identification systems. Countries such as China, Taiwan, South Korea, Vietnam, and Japan play significant roles in global trade. This leads to innovation as reflected by the contribution of Asian countries to the scientific publications. Substantial investment in research and development by Asian governments and companies further supports this innovation [88, 89].
@@
Public funding is reported as the main source, supporting about two out of five publications, indicating strong governmental backing for the logistics sector that operates with low profit margins [90, 91]. The motivation for strong governmental backing is the potential for technological advancements that will benefit national infrastructure and economic growth [92, 93]. Besides this dominant public sector support, industry funding and public-industry collaborations constitute one ninth of the publications, suggesting that while there is some industry involvement, companies largely rely on public sector investments to drive technological development of ILU identification systems. Nevertheless, the presence of pure industry funding (4,76 %) indicates a genuine market demand for companies to invest in research to remain competitive [94, 4].
Figure 3: Ranking of publication outlets for the eligible articles ($n=63$) over the year. The years with zero eligible publications are ommited in this barchart. As this scoping review was prepared in 2025, the data for that year remain incomplete, covering only publications up to August.
@@
While funding sources indicate industry and governmental priorities, the ranking of publication venues offers insight into the field’s academic recognition. To analyse the rankings of journals and conferences we followed the rankings of Scimago [^2] https://www.scimagojr.com/ for journals and CORE [^3] https://www.core.edu.au/icore-portal for conferences. Figure 3 shows that more than half of the articles on automatic ILU identification task are published in unranked outlets, which is not uncommon for applied or niche technical topics [95]. Especially, the early years of DL dominance marked a period of stagnation (2017-2019), where the field struggled to deliver methodological innovation. This trend is changing in the recent years as the percentage of articles published in ranked outlets are increasing, as well as the total number of publications, suggesting that the ILU identification topic is relevant for high-quality research venues. Nevertheless, the field could benefit from a greater emphasis on more novel contributions, increased comparability, and reproducibility of the results. This could be achieved through publicly available benchmark datasets and repositories that showcase the source code and the model weights. By adopting such research standards, ILU identification tasks with CV-based methods could be more relevant for high-ranked publication outlets.